# Expressive Power of Graph Neural Networks for (Mixed-Integer) Quadratic Programs

### v1

@@ 5. Numerical experiments
A study on the expressive power
We train GNNs to fit $\Phi_{\textup{obj}}$ or $\Phi_{\textup{sol}}$ for LCQP or MI-LCQP instances. [^4] Since LCQP and MI-LCQP are linearly constrained, predicting feasibility falls to the case of LP and MILP, which has been numerically investigated in [chen2022representing-lp] and [chen2022representing-milp]. Hence we omit the feasibility experiments here. For both LCQP and MI-LCQP, we randomly generate 100 instances, each of which contains 10 constraints and 50 variables. The generated MI-LCQPs are all unfoldable and MP-tractable with probability one. The optimal solutions and corresponding objective function values are collected using existing solvers. Details on the data generation and training schemes can be found in Appendix E. We train four GNNs with four different embedding sizes and record their relative errors averaged on all instances during training. [^5] The relative error of a GNN $F_{W}$ on a single problem instance $G$ is defined as $\|F_{W}(G)-\Phi(G)\|_{2}/\max(\|\Phi(G)\|_{2},1)$, where $\Phi$ could be either $\Phi_{\textup{obj}}$ or $\Phi_{\textup{sol}}$. The results are reported in Figure 3. We can see that GNNs can fit $\Phi_{\textup{obj}}$ and $\Phi_{\textup{sol}}$ well for both LCQP and MI-LCQP. These results validate Theorems 3.3,3.4,4.7 and 4.9 on a small set of instances. We also observe that a larger embedding size increases the capacity of a GNN, resulting in not only lower final errors but also faster convergence.
@@
Figure 3. Relative errors when training GNNs to fit $\Phi_{\textup{obj}}$ and $\Phi_{\textup{sol}}$ for LCQP (3(a)-3(b)) and MI-LCQP (3(c)-3(d)). GNNs are trained on 100 randomly generated problem instances.
To further validate the theorems on a larger scale, we expand the number of problem instances to 500 and 2,500, and conduct training on the four GNNs along with a larger variant with an embedding size of 1,024. The results are reported in Figure 4. We can observe that GNN can achieve near-zero fitting errors as long as it has a large enough embedding size and thus enough capacity for approximation.
@@
Figure 4. Best relative errors achieved during training for various combinations of embedding sizes and numbers of training samples. We can achieve near-zero errors as long as the GNN is large enough.
A study on the generalization ability
Besides investigating GNNs’ approximation capacity, we also explore their generalization ability. However, the instances used in previous experiments are generic. Here, we generate new LCQP instances by varying only $c$, while keeping other components fixed. Training GNNs with an embedding size of 512 on these instances and testing them on a separate validation set, we observe in Figure 5 that as the number of training instances increases, the generalization gap decreases and validation error improves. With 25,000 training instances, we achieve a satisfactory validation error below $10^{-2}$ for $\Phi_{\mathrm{obj}}$. Results of fitting $\Phi_{\mathrm{sol}}$ can be found in Appendix E.
Figure 5. Train and validation errors when fitting $\Phi_{\mathrm{obj}}$ for LCQP.
6. Conclusion
This paper establishes theoretical foundations for using GNNs to represent the feasibility, optimal objective value, and optimal solution, of LCQPs and MI-LCQPs. In particular, we prove the existence of GNNs that can predict those properties of LCQPs universally well and show with explicit examples that such results are generally not true for MI-LCQPs when integer constraints are introduced. Moreover, we precisely identify subclasses of MI-LCQP problems on which such universal approximation results are still valid. Our findings are also verified numerically.
Let us also comment on the limitations and future directions. Firstly, our universal approximation theorems only show the existence of the GNNs, without discussing the complexity, training, and generalization, which are all important future directions. Secondly, the research conducted in this paper is mainly theoretical and does not cover large-scale experiments. It would be interesting to explore the probability of building machine-learning-based general solvers for large-scale QPs.
@@ Appendix A Proofs for Section 3
In this appendix, we present the proofs for theorems in Section 3. The proofs will based on Weisfeiler-Lehman (WL) test and its separation power to distinguish LCQP problems with different properties.
The Weisfeiler-Lehman (WL) test [weisfeiler1968reduction] is a classical algorithm for the graph isomorphism problem. In particular, it implements color refinement on vertices by applying a hash function on the previous vertex color and aggregation of colors from neighbors, and identifies two graphs as isomorphic if their final color multisets are the same. It is worth noting that WL test may incorrectly identify two non-isomorphic graphs as isomorphic. We slightly modify the standard WL test to fit the structure of LCQP-graphs, see Algorithm 2.
Algorithm 2 The WL test for LCQP-graphs
@@
6:return The multisets containing all colors $\left\{\left\{C_{i}^{L,V}\right\}\right\}_{i=0}^{m},\left\{\left\{C_{j}^{L,W}\right\}\right\}_{j=0}^{n}$.
@@
For LCQP-graphs $G_{\textup{LCQP}},\hat{G}_{\textup{LCQP}}\in\mathcal{G}_{\textup{LCQP}}^{m,n}$, let $\{\{C_{i}^{L,V}\}\}_{i=0}^{m},\{\{C_{j}^{L,W}\}\}_{j=0}^{n}$ and $\{\{\hat{C}_{i}^{L,V}\}\}_{i=0}^{m},\{\{\hat{C}_{j}^{L,W}\}\}_{j=0}^{n}$ the color multisets output by Algorithm 2 on $G_{\textup{LCQP}}$ and $\hat{G}_{\textup{LCQP}}$.
  1. (a)
    We say $G_{\textup{LCQP}}\sim\hat{G}_{\textup{LCQP}}$ if for all $L\in\mathbb{N}$ and all hash functions, $\{\{C_{i}^{L,V}\}\}_{i=0}^{m}=\{\{\hat{C}_{i}^{L,V}\}\}_{i=0}^{m}$ and $\{\{C_{j}^{L,W}\}\}_{j=0}^{n}=\{\{\hat{C}_{j}^{L,W}\}\}_{j=0}^{n}$.
  2. (b)
    We say $G_{\textup{LCQP}}\stackrel{{\scriptstyle W}}{{\sim}}\hat{G}_{\textup{LCQP}}$ if for all $L\in\mathbb{N}$ and all hash functions, $\{\{C_{i}^{L,V}\}\}_{i=0}^{m}=\{\{\hat{C}_{i}^{L,V}\}\}_{i=0}^{m}$ and $C_{j}^{L,W}=\hat{C}_{j}^{L,W},\ \forall~{}j\in\{1,2,\dots,n\}$.
@@
For any $G_{\textup{LCQP}},\hat{G}_{\textup{LCQP}}\in\mathcal{G}_{\textup{LCQP}}^{m,n}$, if $G_{\textup{LCQP}}\sim\hat{G}_{\textup{LCQP}}$, then $\Phi_{\textup{obj}}(G_{\textup{LCQP}})=\Phi_{\textup{obj}}(\hat{G}_{\textup{LCQP}})$.
@@
For any $G_{\textup{LCQP}},\hat{G}_{\textup{LCQP}}\in\mathcal{G}_{\textup{LCQP}}^{m,n}$ that are feasible and bounded, if $G_{\textup{LCQP}}\sim\hat{G}_{\textup{LCQP}}$, then there exists some permutation $\sigma_{W}\in S_{n}$ such that $\Phi_{\textup{sol}}(G_{\textup{LCQP}})=\sigma_{W}(\Phi_{\textup{sol}}(\hat{G}_{\textup{LCQP}}))$. Furthermore, if $G_{\textup{LCQP}}\stackrel{{\scriptstyle W}}{{\sim}}\hat{G}_{\textup{LCQP}}$, then $\Phi_{\textup{sol}}(G_{\textup{LCQP}})=\Phi_{\textup{sol}}(\hat{G}_{\textup{LCQP}})$.
@@
Fixe $x\in\mathbb{R}^{n}$ and consider the problem
@@
Suppose that there are no collisions of hash functions or their linear combinations when applying the WL test to $G_{\textup{LCQP}}$ and $\hat{G}_{\textup{LCQP}}$ and there are no strict color refinements in the $L$-th iteration. Since $G_{\textup{LCQP}}\sim\hat{G}_{\textup{LCQP}}$, after performing some permutation, there exist =$\mathcal{I}=\{I_{1},I_{2},\dots,I_{s}\}$ and $\mathcal{J}=\{J_{1},J_{2},\dots,J_{t}\}$ that are partitions of $\{1,2,\dots,m\}$ and $\{1,2,\dots,n\}$, respectively, such that the followings hold:
@@
$\frac{1}{2}x^{\top}Qx$ (A.1)${\geq}\frac{1}{2}\hat{x}^{\top}Q\hat{x}=\frac{1}{2}\sum_{q,q^{\prime}=1}^{t}\sum_{j\in J_{q}}\sum_{j^{\prime}\in J_{q^{\prime}}}\hat{x}_{j}Q_{jj^{\prime}}\hat{x}_{j^{\prime}}=\frac{1}{2}\sum_{q,q^{\prime}=1}^{t}y_{q}y_{q^{\prime}}\sum_{j^{\prime}\in J_{q^{\prime}}}Q_{jj^{\prime}}$
@@
Under the same setting as in the proof of Theorem A.2, the results can be proved using the same arguments as in the proof of Lemma B.4 and Corollary B.7 in [chen2022representing-lp]. ∎
@@
It is well-known from previous literature that the separation power of GNNs is equivalent to that of WL test and that GNNs can universally approximate any continuous function whose separation is not stronger than that of WL test; see e.g. [chen2022representing-lp, xu2019powerful, azizian2020expressive, geerts2022expressiveness]. We have established in Theorem A.2, Theorem A.3, and Corollary A.5 that the separation power of $\Phi_{\textup{obj}}$ and $\Phi_{\textup{sol}}$ is upper bounded by the WL test (Algorithm 2) that shares the same information aggregation mechanism as the GNNs in $\mathcal{F}_{\textup{LCQP}}$ and $\mathcal{F}_{\textup{LCQP}}^{W}$. Therefore, Theorem 3.3 and Theorem 3.4 can be proved using standard arguments in the previous literature.
@@
Based on Theorem A.2, Theorem 3.3 can be proved following the same lines as in the proof of Theorem 3.4 in [chen2022representing-lp], with trivial modifications to generalize results for LP-graphs to the LCQP setting. ∎
@@
Based on Theorem A.3 and Corollary A.5, Theorem 3.4 can be proved following the same lines as in the proof of Theorem 3.6 in [chen2022representing-lp], with trivial modifications to generalize results for LP-graphs to the LCQP setting. ∎
@@
This section presents the proofs of Proposition 4.2 and Proposition 4.3.

### v2

@@ 5. Numerical experiments
We present numerical experiments in this section. [^3] Codes are available at https://github.com/liujl11git/GNN-QP.
Numerical validation of GNNs’ expressive power
We train GNNs to fit $\Phi_{\textup{obj}}$ or $\Phi_{\textup{sol}}$ for LCQP or MI-LCQP instances. [^4] Since LCQP and MI-LCQP are linearly constrained, predicting feasibility falls to the case of LP and MILP in [chen2022representing-lp, chen2022representing-milp]. Hence we omit the feasibility experiments. For both LCQP and MI-LCQP, we randomly generate 100 instances, each of which contains 10 constraints and 50 variables. The generated MI-LCQPs are all GNN-solvable and GNN-analyzable with probability one. Details on the data generation and training schemes can be found in Appendix G. We train four GNNs with four different embedding sizes and record their relative errors [^5] The relative error of a GNN $F_{W}$ on a single problem instance $G$ is defined as $\|F_{W}(G)-\Phi(G)\|_{2}/\max(\|\Phi(G)\|_{2},1)$, where $\Phi$ could be either $\Phi_{\textup{obj}}$ or $\Phi_{\textup{sol}}$. averaged on all instances during training. The results are reported in Figure 3. We can see that GNNs can fit $\Phi_{\textup{obj}}$ and $\Phi_{\textup{sol}}$ well for both LCQP and MI-LCQP. These results validate Theorems 3.3,3.4,4.6 and 4.7. We also observe that a larger embedding size increases the capacity of a GNN, resulting in not only lower final errors but also faster convergence.
@@
Figure 3. Relative errors when training GNNs to fit $\Phi_{\textup{obj}}$ and $\Phi_{\textup{sol}}$ for LCQP (3(a)-3(b)) and MI-LCQP (3(c)-3(d)) with various embedding sizes.
Validation on a larger scale
To further validate the theorems, we expand the number of problem instances to 500 and 2,500. The results, reported in Figure 4, show that GNN achieves near-zero fitting errors when it has a large enough embedding size and thus enough capacity for approximation, which directly validate Theorems 3.3,3.4,4.6 and 4.7.
@@
Figure 4. Validation on a larger scale. The figures illustrate the relative errors for various combinations of embedding sizes and numbers of training samples. We can achieve near zero errors when the GNN is large enough.
Various types of LCQP
Besides the generic LCQP formulation (2.1), we also extend the numerical experiments to other types of optimization problems, namely portfolio optimization and support vector machine (SVM) following [jung2022learning]. The results are reported in Figure 5. We can observe similar fitting behaviors as those in the generic LCQP experiments where the expressive power of GNNs increase as they become larger, evidenced by the fitting errors decreasing to near zero when the embedding size increases.
(a) Fit $\Phi_{\textup{obj}}$ for portfolio
(b) Fit $\Phi_{\textup{sol}}$ for portfolio
(c) Fit $\Phi_{\textup{obj}}$ for SVM
(d) Fit $\Phi_{\textup{sol}}$ for SVM
Figure 5. Various types of LCQP: portfolio optimization and SVM optimization problems. The figures illustrate the best relative errors achieved during training for various combinations of embedding sizes and numbers of training samples.
Real dataset
To further examine the universal approximation results on more realistic QP problems. For LCQP, we train GNNs on the Maros and Meszaros Problem Set [maros1999repository], which contains 138 challenging convex LCQPs. The results are shown in Figure 6. We observe that while the broad range of numbers of instances in the Maros Meszaros test set caused numerical difficulties for training, GNNs can still be trained to fit the objectives and solutions to some extent. And we can observe similar tendency as in the synthesized experiments that the expressive power increases as the model capacity enlarges when we increase the embedding size. Details are deferred to Appendix G.
(a) Fit $\Phi_{\textup{obj}}$ on Maros Meszaros
(b) Fit $\Phi_{\textup{sol}}$ on Maros Meszaros
(c) Fit $\Phi_{\textup{obj}}$ on QPLib
(d) Fit $\Phi_{\textup{sol}}$ on QPLib
Figure 6. Training errors on real datasets: Convex LCQP (6(a)-6(b)) and MI-LCQP (6(c)-6(d)).
For MI-LCQP, we perform GNN training on 73 GNN-solvable instances from QPLib [furini2019qplib] that provide the optimal solutions and objectives. We train GNNs of embedding sizes 128, 256, and 512 to fit the objectives and solutions. The training errors we achieved are shown in Figures 6(c)-6(d). GNNs can fit the objective values well and demonstrate the ability to fit solutions. The results show the model capacity improves as the model size increases.
Computation complexity
GNNs are superior over QP solvers in terms of running time, especially when we fully exploit parallel computing with GPU acceleration. To show this, we measure the average running time using OSQP [osqp] and a trained GNN with different batch sizes over the 1,000 synthetic LCQP problems generated in the experiment above. We applied OSQP to solve all instances to a relative error of $10^{-3}$, which is slightly less accurate than the trained GNN (with an average relative error of $6.31\times 10^{-4}$). The average solving times and standard deviations are shown in Table 1. The sufficiently accelerated computation validates GNNs’ capacity as a real-time QP solver or fast warm-start, numerically supporting the rationality of our theoretical study of GNNs for QPs.
Table 1. Average solving times (milliseconds) of GNN and OSQP.
Batch Size
- 1 10 100 1,000
OSQP 4.48±3.62 - - - -
GNN - 53.62±16.72 5.37±1.87 0.504±0.142 0.089±0.002
Generalization
Besides investigating GNNs’ expressive capacity, we also explore their generalization ability and observed positive results. As generalization is beyond this work’s scope, the results are provided in Appendix G.
6. Conclusions and future directions
This paper establishes theoretical foundations for using GNNs to represent the feasibility, optimal objective value, and optimal solution of LCQPs and MI-LCQPs. We prove GNNs can universally approximate these properties for LCQPs and show this is in general not true for MI-LCQPs, except for specific subclasses we identify. Future research directions include studying the training dynamics, generalization behavior, and size requirements of GNNs. In particular, the algorithm-unrolling framework by [yang2024efficient] provides a way to estimate GNN sizes, as it explicitly defines the number of parameters per layer, and prior work (e.g., [li2024power]) shows algorithm-unrolling is a structured GNN. These connections suggest that GNN complexity bounds for (MI-)LCQPs may be derived from the algorithm-unrolling literature. Another compelling direction for future work is to unify LP, LCQP, QCQP, and extensions on higher-order polynomial optimization via hypergraph GNNs.
@@ Appendix A Proofs for Section 3
In this appendix, we present the proofs for theorems in Section 3. The proofs will based on Weisfeiler-Lehman (WL) test and its separation power to distinguish LCQP problems with different properties (different feasibility, different optimal objective, or different optimal solution with the smallest $\ell_{2}$ norm).
The Weisfeiler-Lehman (WL) test [weisfeiler1968reduction] is a classical algorithm for the graph isomorphism problem. In particular, it implements color refinement on vertices by applying a hash function on the previous vertex color and aggregation of colors from neighbors, and identifies two graphs as isomorphic if their final color multisets are the same. It is worth noting that WL test may incorrectly identify two non-isomorphic graphs as isomorphic. We slightly modify the standard WL test, see Algorithm 2.
Algorithm 2 The linear-form WL test
@@
6:return The multisets containing all colors $\{\!\!\{C_{i}^{L,V}\}\!\!\}_{i=0}^{m},\{\!\!\{C_{j}^{L,W}\}\!\!\}_{j=0}^{n}$.
Please note that, Algorithm 2 is a special case of Algorithm 1 and, as such, has weaker separation power. However, it is sufficient for distinguishing LCQPs with different key properties. Notably, proving universal approximation under a weaker WL test leads to an even stronger conclusion: if a weaker WL test (or equivalently, a less powerful GNN) can separate LCQPs, it automatically implies that these LCQPs can also be separated under a stronger WL test.
@@
For two LCQP-graphs $G_{\textup{LCQP}},\hat{G}_{\textup{LCQP}}\in\mathcal{G}_{\textup{LCQP}}^{m,n}$, let $\{\!\!\{C_{i}^{L,V}\}\!\!\}_{i=0}^{m},\{\!\!\{C_{j}^{L,W}\}\!\!\}_{j=0}^{n}$ and $\{\!\!\{\hat{C}_{i}^{L,V}\}\!\!\}_{i=0}^{m},\{\!\!\{\hat{C}_{j}^{L,W}\}\!\!\}_{j=0}^{n}$ be color multisets output by Algorithm 2 on $G_{\textup{LCQP}}$ and $\hat{G}_{\textup{LCQP}}$.
  1. (1)
    We say $G_{\textup{LCQP}}\sim\hat{G}_{\textup{LCQP}}$ if $\{\!\!\{C_{i}^{L,V}\}\!\!\}_{i=0}^{m}=\{\!\!\{\hat{C}_{i}^{L,V}\}\!\!\}_{i=0}^{m}$ and $\{\!\!\{C_{j}^{L,W}\}\!\!\}_{j=0}^{n}=\{\!\!\{\hat{C}_{j}^{L,W}\}\!\!\}_{j=0}^{n}$ hold for all $L\in\mathbb{N}$ and all hash functions.
  2. (2)
    We say $G_{\textup{LCQP}}\stackrel{{\scriptstyle W}}{{\sim}}\hat{G}_{\textup{LCQP}}$ if $\{\!\!\{C_{i}^{L,V}\}\!\!\}_{i=0}^{m}=\{\!\!\{\hat{C}_{i}^{L,V}\}\!\!\}_{i=0}^{m}$ and $C_{j}^{L,W}=\hat{C}_{j}^{L,W},\ \forall~j\in\{1,2,\dots,n\}$, for all $L\in\mathbb{N}$ and all hash functions.
@@
For any $G_{\textup{LCQP}},\hat{G}_{\textup{LCQP}}\in\mathcal{G}_{\textup{LCQP}}^{m,n}$ with $Q,\hat{Q}\succeq 0$, if $G_{\textup{LCQP}}\sim\hat{G}_{\textup{LCQP}}$, then $\Phi_{\textup{obj}}(G_{\textup{LCQP}})=\Phi_{\textup{obj}}(\hat{G}_{\textup{LCQP}})$.
@@
For any $G_{\textup{LCQP}},\hat{G}_{\textup{LCQP}}\in\mathcal{G}_{\textup{LCQP}}^{m,n}$ with $Q,\hat{Q}\succeq 0$ that are feasible and bounded, if $G_{\textup{LCQP}}\sim\hat{G}_{\textup{LCQP}}$, then there exists some permutation $\sigma_{W}\in S_{n}$ such that $\Phi_{\textup{sol}}(G_{\textup{LCQP}})=\sigma_{W}(\Phi_{\textup{sol}}(\hat{G}_{\textup{LCQP}}))$. Furthermore, if $G_{\textup{LCQP}}\stackrel{{\scriptstyle W}}{{\sim}}\hat{G}_{\textup{LCQP}}$, then $\Phi_{\textup{sol}}(G_{\textup{LCQP}})=\Phi_{\textup{sol}}(\hat{G}_{\textup{LCQP}})$.
@@
Fix $x\in\mathbb{R}^{n}$ and consider the problem
@@
Suppose that there are no collisions of hash functions or their linear combinations when applying the WL test to $G_{\textup{LCQP}}$ and $\hat{G}_{\textup{LCQP}}$ and there are no strict color refinements in the $L$-th iteration. Since $G_{\textup{LCQP}}\sim\hat{G}_{\textup{LCQP}}$, after performing some permutation, there exist $\mathcal{I}=\{I_{1},I_{2},\dots,I_{s}\}$ and $\mathcal{J}=\{J_{1},J_{2},\dots,J_{t}\}$ that are partitions of $\{1,2,\dots,m\}$ and $\{1,2,\dots,n\}$, respectively, such that the followings hold:
@@
[…]
@@
Under the same setting as in the proof of Theorem A.2, the results can be proved. We present the proof here.
Let $x\in\mathbb{R}^{n}$ be the optimal solution to (2.1) with the smallest $\ell_{2}$-norm, and let $\hat{x}\in\mathbb{R}^{n}$ be defined as in the proof of Theorem A.2. By the arguments in the proof of Theorem A.2, $\hat{x}$ is an optimal solution to (A.3). In particular, $\hat{x}$ is also an optimal solution to (2.1) since one can set $(\hat{A},\hat{b},\hat{c},\hat{Q},\hat{l},\hat{u},\hat{\circ})=(A,b,c,Q,l,u,\circ)$. Therefore, by the minimality of $\|x\|^{2}$, we have that
$\|x\|^{2}\leq\|\hat{x}\|^{2}=\sum_{q=1}^{t}\sum_{j\in J_{q}}\hat{x}_{j}^{2}=\sum_{q=1}^{t}|J_{q}|\left(\frac{1}{|J_{q}|}\sum_{j\in J_{q}}x_{j}\right)^{2}\leq\sum_{q=1}^{t}\sum_{j\in J_{q}}x_{j}^{2}=\|x\|^{2},$
which implies that $x_{j}$ is a constant in $j\in J_{q}$ and $x=\hat{x}$. Thus, $x$ is also an optimal solution to (A.3).
Let $x^{\prime}\in\mathbb{R}^{n}$ be the optimal solution to (A.3) with the smallest $\ell_{2}$-norm. Then $\|x^{\prime}\|\leq\|\hat{x}\|=\|x\|$ and the reverse direction $\|x\|\leq\|x^{\prime}\|$ is also true, which implies that $\|x\|=\|x^{\prime}\|$. Therefore, we have $x=x^{\prime}$ by the uniqueness of the optimal solution with the smallest $\ell_{2}$-norm.
Noticing that the above arguments are made after permuting vertices in $V$ and $W$, we can conclude that $\Phi_{\textup{sol}}(G_{\textup{LCQP}})=\sigma_{W}(\Phi_{\textup{sol}}(\hat{G}_{\textup{LCQP}}))$ for some $\sigma_{W}\in S_{n}$. Additionally, if $G_{\textup{LCQP}}\stackrel{{\scriptstyle W}}{{\sim}}\hat{G}_{\textup{LCQP}}$, then there is no need to perform the permutation on $W$ and we have $\Phi_{\textup{sol}}(G_{\textup{LCQP}})=\Phi_{\textup{sol}}(\hat{G}_{\textup{LCQP}})$. ∎
@@
It is well-known from previous literature that the separation power of GNNs is equivalent to that of WL test and that GNNs can universally approximate any continuous function whose separation is not stronger than that of WL test; see e.g. [chen2022representing-lp, xu2019powerful, azizian2020expressive, geerts2022expressiveness]. We have established in Theorem A.2, Theorem A.3, and Corollary A.5 that the separation power of $\Phi_{\textup{obj}}$ and $\Phi_{\textup{sol}}$ is upper bounded by the WL test (Algorithm 2) that shares the same information aggregation mechanism as the GNNs in $\mathcal{F}_{\textup{LCQP}}$ and $\mathcal{F}_{\textup{LCQP}}^{W}$. Theorem 3.3 and Theorem 3.4 will be proved following this idea.
@@
Theorem 3.3 can be proved based on Theorem A.2. The separation power of GNNs is equivalent to that of the WL test, i.e., for any $G_{\textup{LCQP}},\hat{G}_{\textup{LCQP}}\in\mathcal{G}_{\textup{LCQP}}^{m,n}$ with $Q,\hat{Q}\succeq 0$,
(A.4) $G_{\textup{LCQP}}\sim\hat{G}_{\textup{LCQP}}\Longleftrightarrow F(G_{\textup{LCQP}})=F(\hat{G}_{\textup{LCQP}}),\ \forall~F\in\mathcal{F}_{\textup{LCQP}},$
which combined with Theorem A.2 leads to that
(A.5) $F(G_{\textup{LCQP}})=F(\hat{G}_{\textup{LCQP}}),\ \forall~F\in\mathcal{F}_{\textup{LCQP}}\implies\Phi_{\textup{obj}}(G_{\textup{LCQP}})=\Phi_{\textup{obj}}(\hat{G}_{\textup{LCQP}}),$
indicating that the separation power of $\mathcal{F}_{\textup{LCQP}}$ is upper bounded by that of $\Phi_{\textup{obj}}$.
The indicator function $\mathbb{I}_{\Phi_{\textup{obj}}(\cdot)\in\mathbb{R}}:\mathcal{G}_{\textup{LCQP}}^{m,n}\to\{0,1\}\subset\mathbb{R}$ is measurable, i.e., $\Phi_{\textup{obj}}^{-1}(0)$ and $\Phi_{\textup{obj}}^{-1}(0)$ are both Lebesgue measurable, and hence by Lusin’s theorem, there exists a compact and permutation-invariant subspace $X\subset\mathcal{G}_{\textup{LCQP}}^{m,n}$ such that $\mathbb{P}[\mathcal{G}_{\textup{LCQP}}^{m,n}\backslash X]<\epsilon$ and that $\mathbb{I}_{\Phi_{\textup{obj}}(\cdot)\in\mathbb{R}}$ restricted on $X$ is continuous. Therefore, by the Stone-Weierstrass theorem and (A.5), we have that there exists $F_{1}\in\mathcal{F}_{\textup{LCQP}}$ satisfying
$\sup_{G_{\textup{LCQP}}\in X}\left|F_{1}(G_{\textup{LCQP}})-\mathbb{I}_{\Phi_{\textup{obj}}(G_{\textup{LCQP}})\in\mathbb{R}}\right|<\frac{1}{2}$
Therefore, it holds that
[…]
which proves (3.1). Additionally, (3.2) can be proved by applying similar arguments to $\Phi_{\textup{obj}}:\Phi_{\textup{obj}}^{-1}(\mathbb{R})\to\mathbb{R}$, where $\Phi_{\textup{obj}}^{-1}(\mathbb{R})\subset\mathcal{G}_{\textup{LCQP}}^{m,n}$ is the collection of feasible and bounded $G_{\textup{LCQP}}\in\mathcal{G}_{\textup{LCQP}}^{m,n}$. ∎
@@
Theorem 3.4 can be proved based on Theorem A.3 and Corollary A.5. In addition to (A.4), it can be proved that the separation powers of GNNs and the WL test are equivalent in the following sense:
  • For any $G_{\textup{LCQP}},\hat{G}_{\textup{LCQP}}\in\mathcal{G}_{\textup{LCQP}}^{m,n}$, $G_{\textup{LCQP}}\stackrel{{\scriptstyle W}}{{\sim}}\hat{G}_{\textup{LCQP}}$ if and only if $F_{W}(G_{\textup{LCQP}})=F_{W}(\hat{G}_{\textup{LCQP}})$ for all $F_{W}\in\mathcal{F}_{\textup{LCQP}}^{W}$.
  • For any $G_{\textup{LCQP}}\in\mathcal{G}_{\textup{LCQP}}^{m,n}$ and any $j,j^{\prime}\in W$, $C_{j}^{L,W}=C_{j^{\prime}}^{L,W}$ for any $L\in\mathbb{N}$ and any hash function if and only if $F_{W}(G_{\textup{LCQP}})_{j}=F_{W}(G_{\textup{LCQP}})_{j^{\prime}}$ for all $F_{W}\in\mathcal{F}_{\textup{LCQP}}^{W}$.
Therefore, with Theorem A.3 and Corollary A.5, the separation power of GNNs is upper bounded by that of $\Phi_{\textup{sol}}$ in the following sense that for any $G_{\textup{LCQP}},\hat{G}_{\textup{LCQP}}\in\mathcal{G}_{\textup{LCQP}}^{m,n}$ with $Q,\hat{Q}\succeq 0$ and any $j,j^{\prime}\in W$,
  • $F(G_{\textup{LCQP}})=F(\hat{G}_{\textup{LCQP}}),\ \forall~F\in\mathcal{F}_{\textup{LCQP}}$ implies $\Phi_{\textup{sol}}(G_{\textup{LCQP}})=\sigma_{W}(\Phi_{\textup{sol}}(\hat{G}_{\textup{LCQP}}))$ for some $\sigma_{W}\in S_{n}$.
  • $F_{W}(G_{\textup{LCQP}})=F_{W}(\hat{G}_{\textup{LCQP}}),\ \forall~F_{W}\in\mathcal{F}_{\textup{LCQP}}^{W}$ implies $\Phi_{\textup{sol}}(G_{\textup{LCQP}})=\Phi_{\textup{sol}}(\hat{G}_{\textup{LCQP}})$.
  • $F_{W}(G_{\textup{LCQP}})_{j}=F_{W}(G_{\textup{LCQP}})_{j^{\prime}},\ \forall~F_{W}\in\mathcal{F}_{\textup{LCQP}}^{W}$ implies $\Phi_{\textup{sol}}(G_{\textup{LCQP}})_{j}=\Phi_{\textup{sol}}(G_{\textup{LCQP}})_{j^{\prime}}$.
The optimal solution mapping $\Phi_{\textup{sol}}:\Phi_{\textup{obj}}^{-1}(\mathbb{R})\to\mathbb{R}^{n}$ is measurable, i.e., $\Phi_{\textup{sol}}^{-1}(A)$ is Lebesgue measurable for any Borel measurable $A\subset\mathbb{R}^{n}$, and hence by Lusin’s theorem, there exists a compact and permutation-invariant subspace $X\subset\Phi_{\textup{obj}}^{-1}(\mathbb{R})$ such that $\mathbb{P}[\Phi_{\textup{obj}}^{-1}(\mathbb{R})\backslash X]<\epsilon$ and that $\Phi_{\textup{sol}}$ restricted on $X$ is continuous. Therefore, applying the generalized Stone-Weierstrass theorem for equivariant functions [azizian2020expressive]*Theorem 22, we know that there exists $F_{W}\in\mathcal{F}_{\textup{LCQP}}^{W}$ satisfying
$\sup_{G_{\textup{LCQP}}\in X}\left\|F_{W}(G_{\textup{LCQP}})-\Phi_{\textup{sol}}(G_{\textup{LCQP}})\right\|<\delta.$
Therefore, it holds that
[…]
which completes the proof. ∎
@@
The proof of Proposisition 4.1 is directly from [chen2022representing-milp] since adding a quadratic term in the objective function of an MILP problem does not change the feasible region. However, Proposisitions 4.2 and 4.3 are not covered in [chen2022representing-milp] and we present their proofs here.